Substitution method:
3x-6y=-2
2x-3y=1
Post steps used to solve.
Thank you.

Answer by Alan3354(69443)

Substitution method:
3x-6y=-2
2x-3y=1 x2 --> 4x-6y = 2 --> 6y = 4x - 2
-------
3x-6y=-2
Sub for 6y
3x - (4x-2) = -2
-x + 2 = -2
x = 4
---------
2x-3y=1
2*4 - 3y = 1
-3y = -7
y = 7/3